AB; crystallizes in a body centred cubic lattice with edge length ‘a equal to 387pm. The distance between two oppositely charged ions
in the lattice is :
(a) 335 pm
(b) 250 pm
(c) 200 pm
(d) 300 pm

Grade:9

2 Answers

Vikas TU
6 years ago
The distance b/w them would be ;
2r = root(3)*a/2
2r => 387*root3/2 => 335 pm approx.
This would be the distance b/w the two opposite charges.
Athira
5 years ago
The distance between opposite charges would be :
r=√3*a/4 here the edge length ,a is given by 387pm 
Implies r = √3 *387 /4 
We get 335 approximately ..